Silicon dioxide will react with aluminum to form silicon and aluminum oxide. Write the balanced equation for the reaction and report the reaction type. Calculate the mass of aluminum oxide produced if 7.004 g of aluminum reacts with 10.83 g of silicon dioxide.
Answer #1

1)

Balanced chemical equation is:

3 SiO2 + 4 Al ---> 2 Al2O3 + 3 Si

2)

Molar mass of SiO2,

MM = 1*MM(Si) + 2*MM(O)

= 1*28.09 + 2*16.0

= 60.09 g/mol

mass(SiO2)= 10.83 g

use:

number of mol of SiO2,

n = mass of SiO2/molar mass of SiO2

=(10.83 g)/(60.09 g/mol)

= 0.1802 mol

Molar mass of Al = 26.98 g/mol

mass(Al)= 7.004 g

use:

number of mol of Al,

n = mass of Al/molar mass of Al

=(7.004 g)/(26.98 g/mol)

= 0.2596 mol

Balanced chemical equation is:

3 SiO2 + 4 Al ---> 2 Al2O3 + 3 Si

3 mol of SiO2 reacts with 4 mol of Al

for 0.1802 mol of SiO2, 0.2403 mol of Al is required

But we have 0.2596 mol of Al

so, SiO2 is limiting reagent

we will use SiO2 in further calculation

Molar mass of Al2O3,

MM = 2*MM(Al) + 3*MM(O)

= 2*26.98 + 3*16.0

= 101.96 g/mol

According to balanced equation

mol of Al2O3 formed = (2/3)* moles of SiO2

= (2/3)*0.1802

= 0.1202 mol

use:

mass of Al2O3 = number of mol * molar mass

= 0.1202* 101.96

= 12.25 g

Answer: 12.2 g